OpenStudy (anonymous):

Let = –5x + 3 and g(x) = 6x – 2. Find f • g and its domain. (1 point) –30x2 + 28x – 6; all real numbers except x = –15x2 – 8x – 12; all real numbers except x = –30x2 + 28x – 6; all real numbers –15x2 – 8x – 12; all real numbers

OpenStudy (anonymous):

\[f \cdot g\] means just multiply the given functions..

OpenStudy (anonymous):

So just do this: \[f \cdot g = (-5x + 3)(6x-2)\]

OpenStudy (anonymous):

the combine like terms?

OpenStudy (anonymous):

You know about FOIL..??

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

-30x+10+18-6?

OpenStudy (anonymous):

It will be: -30x^2 + 10x + 18x - 6 Now combine like terms..

OpenStudy (anonymous):

–30x2 + 28x – 6;

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

does x=1/3?

OpenStudy (anonymous):

Here, you have to find domain, that is value of x for which -30x^2 + 28x - 6 will become undefined.. So, is there any value??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Yep.. So, domain is All Reals..